OpenStudy (rane):

how to use the gradient from the velocity vs tan angle graph to find the value of g ( gravity) ?????

OpenStudy (rane):

@Mashy

OpenStudy (anonymous):

so magnitude of velocity vs Tangent of the angle, that the velocity makes with x axis or something?

OpenStudy (anonymous):

i have never worked with such graphs :D

OpenStudy (rane):

so u don't know how to work out g ( gravity ) form the gradient ?

OpenStudy (anonymous):

I am not sure i understand the question your graph.. tell me what does it have along y axis nd along x axis?

OpenStudy (rane):

y is the velocity and on the x-axis is the Tangent of the angle,

OpenStudy (anonymous):

angle between what?!

OpenStudy (anonymous):

Besides what sort of graph is that.. i have nEver heard of something like that before.. why do you even use that o.O

OpenStudy (rane):



OpenStudy (anonymous):

Ohhww.. are you talking about circular motion!?

OpenStudy (anonymous):

i really have no clue.. u better ask for someone else :P

OpenStudy (rane):

ok

OpenStudy (rane):

thats derivative

OpenStudy (anonymous):

This is your answer .... @RANE

OpenStudy (rane):

the formula for the gravity is \[g = \frac{ v ^{2} }{ r \tan \theta } \]

OpenStudy (rane):

how is the value of the gradient can be used in this equation ?

OpenStudy (anonymous):

Ok Hold ... I will tell you ....

OpenStudy (rane):

ok

OpenStudy (anonymous):

@RANE I can't find the answer ..... Your question is too critical ....

OpenStudy (rane):

oh

OpenStudy (anonymous):

May be you can find your answer in Google ....

OpenStudy (rane):

its not there i have tried it

OpenStudy (anonymous):

Then I think so there is no answer of this question ....

OpenStudy (rane):

there has to be its a lab report

OpenStudy (anonymous):

Then I prefer you that contact any lab .... :P

OpenStudy (rane):

lol, i wish

OpenStudy (shadowlegendx):

I'm sorry but I'm not good with physics and its late for me and I'm tired :P Not many physics masters are on atm but these are some :3 @Callisto @whpalmer4 @hoblos @wolfe8

OpenStudy (mrnood):

If your equation is correct then Tan theta = v^2/rg If you plot tan theta against V^2/r then you should have a straight line - the gradient will be 1/g (Note : r = L cosTheta) I know that his is not what you asked but - since the graph you describe is a curve it does not have a general value for gradient.

OpenStudy (rane):

i didn't say it a curve or anything like that but ust to tell you its straight line

OpenStudy (anonymous):

@RANE @MrNood @ShadowLegendX I want to say that there is no answer ..... So,please Don't waste your Time ... :P

OpenStudy (rane):

@arslanfarooka some of my class mates have done it there is an an

OpenStudy (rane):

answer*

OpenStudy (anonymous):

I do my best But I can't find the answer .... I will Contact my professor Then I shall tell you the proper answer .... Is that ok ...

OpenStudy (rane):

no its fine thanks for your concern

OpenStudy (mrnood):

@arslanfarooka On what basis do you say there is not an answer? If the data exists and the formula is correct then even if there is not an analytical solution there must be an empirical one...

OpenStudy (anonymous):

@MrNood Bro In this question there is no data exists and have no formula then what kind of method that we used to find the proper answer .... tell me hannn ... :@

OpenStudy (rane):

ok guys relax pls,we r not here to discuss or argue weather these is formula or not but wht is the way to solve it

OpenStudy (anonymous):

@RANE The only way to solve this that you contact your professor or your seniors .... Got it .... @RANE If you find the answer then please Inboxed me ... :)

OpenStudy (rane):

ok i will

OpenStudy (mrnood):

OK - I have looked at the diagram - I think I see what this is - You are measuring the speed of circular motion over different length of string If oyu resolve the tension (T) in the string: Vertically Tcos thets = mg Horizontally Tsintheta = mv^2/r (centripetal force) Divide the 2 equations and you get Tan Theta = v^2/rg This is how your equation is derived , and shows that the equation is correct. SO gTan theta = v^2/r

OpenStudy (mrnood):

Now - the eqation of a STRAIGHT line is y = mx +c where m is the gradient on your graph y is V and x is tan theta For circular motion v = r w (w= omega= angular velocity) so v^2 /r = w^2 so gtan theta= w^2 In order to make this a straight line you need to plot w^2 in Y and tan theta in x I cannot see how v against tan theta can be a straight line

OpenStudy (mrnood):

Correction v = rw v^2= r^2 w^2 therefore v^/r = rw^2

OpenStudy (anonymous):

@MrNood ok fine bro u won :)

OpenStudy (rane):

thats not th right way but thanks anyway u told me a hard and long way , thank again

OpenStudy (anonymous):

@RANE r u find the answer ...???

OpenStudy (rane):

yh i have figured it out

OpenStudy (mrnood):

@rane care to post it for our education? I'd like to understand the solution....

OpenStudy (anonymous):

@rane send me answer in inbox